# Hofstadter-Toda spectral duality and quantum groups

### v1

@@ 1. Preliminaries on mathematical structures and models
where we used again the fact that the exponential of the momentum corresponds to the translation operator. Setting $\nu_{x}=0$, the left-hand side above is an operator of $\ell^{2}(\mathbb{Z})$ known as the almost Mathieu operator, featured in Barry Simon’s fifteen problems about Schrödinger operators “for the twenty-first century” [Sim00]. This operator comes with a parameter $R>0$ which multiplies the cosine function in Eq. (12).
Let us point out that the measure-theoretic properties of the spectrum of the almost Mathieu operator crucially depend on $R$. When $\alpha$ is irrational, the spectrum is a Cantor set of Lebesgue measure $\lvert 4-4R\rvert$. Hence, the case $R=1$ has spectrum of measure zero, and it can be shown to be surely purely singular continuous spectrum. The case $R>1$ is notable for having almost surely pure point spectrum and exhibiting Anderson localization (see [AA80, AJ06, Avi08]).
@@
It will be convenient to set $T_{x}=\mathrm{e}^{\mathrm{i}\nu_{x}}V,T_{y}=\mathrm{e}^{\mathrm{i}\nu_{y}}U$. The group spanned by $T_{x},T_{y}$ is referred to as the group of magnetic translations [Zak64].
Our goal is now to find an expression for ${\hat{H}}$ in terms of quantum group elements $E,F$, in analogy with the Toda model, in the spirit of [WZ94]. Towards this goal, we define a finite-dimensional representation $\varphi$ of $\mathcal{U}_{q}(\mathfrak{sl}_{2})$ in terms of shift and clock operators. The notation suggests that this should be a variant of the representation $\phi$ introduced above in the Toda model setting. Choosing $t=q^{-1}$ in Proposition 1.3 and relabeling the Clifford generators as above, we define $\varphi$ as follows:
@@
2. Characteristic polynomial and spectral duality
As already mentioned, Hatsuda, Katsura, and Tachikawa [HKT16] found a relationship between the characteristic polynomial of the Hofstadter Hamiltonian $H$, the self-similarity relation defining the fractal properties of the Hofstadter butterfly, and the spectral transformation between the Toda Hamiltonian $H_{\text{Toda}}$ and its modular dual.
Let us denote the polynomial induced by the modular duality between the Toda Hamiltonian $H_{\text{Toda}}$ and its modular dual defined in [HKT16] by $f=f_{P/Q}(E)$, where $E$ is the energy. We are going to prove that $f$ is the characteristic polynomial $f(E)=\det({\hat{H}}-E)$ of the finite-dimensional Hofstadter Hamiltonian (note that ${\hat{H}}$ depends on $P/Q$, although we suppressed this dependence from the notation).
@@ 2. Characteristic polynomial and spectral duality
Let us briefly introduce some elements of spectral duality for tridiagonal matrices (see [Mol97] for more details). The Harper equation can be written in matrix form:
@@
When $\alpha=P/Q$ is a rational number, the energy spectrum of the Hofstadter model splits into $Q$ energy bands. The eigenvalues of ${\hat{H}}$ give us a point in each band, and the rest of the band is obtained by parametrizing over a torus, which coincide with the Brillouin zone, whose coordinates are given by the vector $\vec{\nu}=(\nu_{x},\nu_{y})$. The way $f$ depends on $\vec{\nu}$ is well-understood [Cha65, Koh89]: it is a simple translation as shown by the Chambers relation,
@@
Note that Eq. (24) is in fact a proof of the Chambers relation for the $x$-coordinate. We will see how to obtain the $y$-dependence later after using the representation theory of the quantum group. We will refer to the points $\vec{\nu}=(0,0)$, $\vec{\nu}=(\textstyle\frac{\pi}{2Q},\frac{\pi}{2Q})$, and $\vec{\nu}=(\textstyle\frac{\pi}{Q},\frac{\pi}{Q})$ as the center, mid-band, and corner points, and restrict ourselves to the reduced Brillouin zone $\nu_{x},\nu_{y}\in[0,2\pi/Q]$, except when explicitly noted. This can be done without loss of generality, since the Hamiltonian is periodic in $\nu_{x},\nu_{y}$ with a period $2\pi/Q$ up to unitary transformations [MCO15].
It is well-known that the characteristic polynomial in $f(E,\nu_{x},\nu_{y})$ (of order $Q$) contains only terms with powers $E^{Q-2i}$. This is because odd powers in $E$ necessarily contains terms in $\cos{(2\pi n\alpha-\nu_{y})}$ ), which cancel each other when summed over $n=0,Q-1$ at the mid-band point $\nu_{y}=2\pi/Q$. Consequently, $f(E,\nu_{x},\nu_{y})$ is an even function of $E$ and the spectrum is symmetric under $E\mapsto-E$ if $Q$ is even, while $f(E,\nu_{x},\nu_{y})$ is an odd function and the spectrum at the mid-band point contains $E=0$ if $Q$ is odd, since $f(0,{\pi}/{2Q},{\pi}/{2Q})=0$. It is also well-known that $f(0,{\pi}/{2Q},{\pi}/{2Q})=4(-1)^{Q/2}$ if $Q$ is even, which mandates that $f(0,0,0)=0$ if $Q/2$ is even and $f(0,\pi/Q,\pi/Q)=0$ if $Q/2$ is odd. Consequently, the spectrum contains $E=0$ at the center point if $Q$ is doubly even, at the corner point if $Q$ is singly even, and at the mid-band point if $Q$ is odd, respectively. In particular, the zeros $E=0$ in the spectra when $Q$ is even are doubly degenerate and form Dirac cones with a linear dependence on the momentum [WZ89].
Let us consider the Hamiltonian $H_{\text{Toda}}$ and its modular dual. Since these commute, the authors in [HKT16] argue that, by simultaneous diagonalization, we arrive at a pair of difference equations which can be iterated, revealing a tracial relationship of the form
@@
Since $T_{j}$ and $\tilde{T}_{j}$ are defined by shifting along the imaginary axis, at $x=0$ the hyperbolic cosine function becomes a simple cosine function. Then the left-hand side of Eq. (26), evaluated at $x=0$, ought to be the trace of the transfer matrix from Eq. (23) (there is a slight difference due to the convention in [HKT16] that $[x,p]=2\pi\mathrm{i}\alpha$, but after renormalizing accordingly, the identification with the matrix $T(E)$ in Eq. (23) holds). Then by Theorem 2.1 and Eq. (24), the $x$-dependent part in Eq. (26) corresponds to the $\nu_{y}$-term in the Chambers relation.
In particular, since the right-hand side of Eq. (26) is defined via modular duality (swapping $P$ with $Q$), the correction in the definition of $\tilde{T}$ is such that the $x$-dependent terms are always equal for any $P$ and $Q$. Since the polynomial (denoted $P$ in [HKT16]) is defined by effectively setting to zero the $x$-dependent part, we have the identification $P_{\alpha}(E)=(-1)^{Q}f_{P/Q}(E,\pi/2Q,\pi/2Q)$. ∎
Remark 2.3.
In [HKT16] the authors define a polynomial $P$ including the anisotropy parameter $R$, in which case the relevant polynomial is $P_{\alpha}(E)+2R^{Q}$ (note we are adding $2R^{Q}$ rather than $2$).
3. Formula for the characteristic polynomial
We now work towards finding a formula for $f(E)=\det({\hat{H}}-E)$. We can exploit the gauge invariance of $H_{\text{Hof}}$ and choose a different form for the vector potential. Following [WZ94], we can choose $\vec{A}=\alpha/2\cdot(-x-y,\,x+y+1)$. In this gauge, the mid-band point corresponds to $\vec{\nu}=(\pi/2,\pi/2)$. Taking this into account, we can rewrite the representation $\varphi$ from Eq. (21) in terms of magnetic translations:
@@ 3. Formula for the characteristic polynomial
As a preliminary step, we are interested in the representation $\varphi=\zeta_{0}$ described in [WZ94], given as
@@
From the representation theory of $\mathcal{U}_{q}(\mathfrak{sl}_{2})$, when $q^{2}$ is a primitive root of unity, we have essentially three families of representations. The first two are low-dimensional, and hence they must be discarded in favor of the last one, which we denote $Z_{a,b}(\lambda)$. This family is topologically parametrized by a $3$-dimensional complex space. However, our system naturally comes with a single parameter $R$, which we have seen in Section 1 together with the almost Mathieu operator.
@@
By gauge invariance of $H_{\text{Hof}}$, we can compute $\det({\hat{H}}-E)$ when the vector potential is chosen as $\vec{A}=\alpha/2\cdot(-x-y,\,x+y+1)$. In this case, the results in [WZ94] imply that ${\hat{H}}$ is given as prescribed by Eq. (27) with $\varphi=\zeta_{0}$. According to the classification of the quantum group representations, $\zeta_{0}$ must be equivalent to $Z_{\lambda}$ for some $\lambda$. Indeed, a direct computation (by conjugation with a diagonal matrix) shows that $\lambda=q^{-1}$ (note that, if $\Lambda$ is such diagonal matrix, the system resulting from $\Lambda\zeta_{0}(F)\Lambda^{-1}=\zeta(F)$ is overdetermined). Equivalent representations are related through similar matrices, i.e., which are equivalent up to a similarity (but not necessarily unitary) transformation, and it is well-known that similarity preserves the characteristic polynomial. ∎
Remark 3.3.
Note that ${\hat{H}}$ is hermitian, whereas ${\hat{H}}^{\prime}$ is not. However, the matrices ${\hat{H}}$ and ${\hat{H}}^{\prime}$ are equivalent up to a similarity transformation, and therefore isospectral with real spectra: Indeed, ${\hat{H}}^{\prime}$ is pseudo-hermitian, following the definitions introduced by Mostafazadeh [Mos02]. Hence, the corresponding Hamiltonian $H^{\prime}$ is an example of a non-Hermitian Hamiltonian with real spectrum which is equivalent up to a similarity transformation to a more conventional hermitian Hamiltonian [Mos02, Fer15].
Remark 3.4.
We see from Definition 1.5 that (at least for $\nu_{y}=0$) the diagonal of ${\hat{H}}$ is given by the Chebyshev polynomials of the first kind, in symbols $\theta_{k}=2T_{k-1}(\cos(2\pi\alpha))$. On the other hand, the Hamiltonian ${\hat{H}}^{\prime}$ in the new “quantum group-adjusted” gauge [WZ94] shows the Chebyshev polynomials of the second kind on the upper diagonal, that is, ${\hat{H}}^{\prime}_{k,k+1}=U_{k-1}^{2}(\sin(\pi\alpha))$.
Remark 3.5.
@@
where in the last equality we reindexed starting from $1$ rather than 00. When $i=0$ we get $x^{N}$ as expected. For the other indices we can sum the two summations and use Lemma 3.6 to obtain the result. Analogously, the case where $N$ is even is carried out using the corresponding identities in Lemma 3.6. ∎
@@
Combining the previous theorem with Proposition 3.2, we get the main result on the polynomial $f$. Recall that $\alpha=P/Q$ is a rational number in reduced form.
Theorem 3.9.
@@
Remark 3.10.
Let us note that the formula from Theorem 3.7 makes at least two aspects clear: if we multiply $y$ by a quantity $x$ and the $b_{i}$’s by the inverse $x^{-1}$, the determinant of $A$ is unaffected. Moreover, when $N$ is odd, the determinant goes to zero if we set $x=0$. We thus recover the well-known result mentioned earlier: the spectrum of the Hofstadter model at the mid-band point always contains 00 when $Q$ is odd.
Remark 3.11.
@@
4. Chambers relation and Chebyshev polynomials
@@
Firstly, we know from Section 1 that these Hamiltonians have the form $(q-q^{-1})\phi(E+F)$ for a suitable representation. Secondly, we obtained an expression of the form ${\hat{H}}^{\prime}=\mathrm{i}(q-q^{-1})\zeta(-E+F)=(q-q^{-1})\zeta(\mathrm{i}^{-1}E+\mathrm{i}F)$. The natural guess away from the mid-band point is to set
@@
We can now use Corollary 3.8 to quantify the corner contributions, leading us to the Chambers relation in the coordinates $\kappa_{\pm}$. Let us consider the contribution of one corner, multiplied by $z^{Q}+z^{-Q}$,
@@
We claim that Eq. (32) prescribes the energy dependence on $\vec{\nu}$ (see Theorem 4.2 and its proof for the precise statement).
To prove the claim, inspired by Remark 3.4, we could proceed by induction on $Q$ and use the recurrence relations of the Chebyshev polynomials in the proof. Although this seems viable, we will follow a more direct approach.
Lemma 4.1.
@@
can be proved by standard techniques, see for example [Kre94, Section 4]. We will not go into the details here and rather focus on the particular form of the “offset” function $g$.
@@
From Eq. (25) and Proposition 3.2 we know we only need to focus on the function $g$. From Eq. (32) and Lemma 4.1, we obtain the corner contribution as follows:
@@
Let us end this section with a small remark about the gauge associated with ${\hat{H}}^{\prime}$ (similar considerations hold for ${\hat{H}}_{c}^{\prime}$). As noted in [WZ94], the Harper equation at the mid-band point in this setting becomes a difference functional equation:
@@
Recall that the form of the vector potential is chosen to be $\vec{A}=\alpha/2\cdot(-x-y,\,x+y+1)$. Hence a convenient coordinate for the Bloch wave function is $l=n+m$. The Harper equation for $\psi=(\psi_{l})$ is obtained from the equation above by setting $z=q^{l}$ and $\psi_{l}=\Psi(q^{l})$. Notice that $l$ ranges in $1,\dots,2Q$, presenting a “doubling” of the period in comparison with the Harper equation in Section 1. Clearly, this is also reflected in the entries of ${\hat{H}}^{\prime}$ where $\pi\alpha$ (rather than $2\pi\alpha$) appears.
@@
More precisely, using Theorems 3.7 & 3.9, we have that $b_{j}=\sin^{2}(j\pi\alpha)/\sin^{2}(\pi\alpha)$, so that $b_{Q}=0$, and $b_{Q+j}=b_{j}$. Since the variables $b_{j}$ enter the formula only through polynomials which are a modification of the elementary symmetric polynomials, these are computed by expanding the linear factorization of the monic polynomial $\lambda^{2}\prod_{1}^{Q-1}(\lambda-b_{i})^{2}$.

### v2

@@ I Preliminaries on mathematical structures and models
where we used again the fact that the exponential of the momentum corresponds to the translation operator. Setting $\nu_{x}=0$, the left-hand side above is an operator of $\ell^{2}(\mathbb{Z})$ known as the almost Mathieu operator, featured in Barry Simon’s fifteen problems about Schrödinger operators “for the twenty-first century”Simon (2000). This operator comes with a parameter $R>0$ which multiplies the cosine function in Eq. (12).
Let us point out that the measure-theoretic properties of the spectrum of the almost Mathieu operator crucially depend on $R$. When $\alpha$ is irrational, the spectrum is a Cantor set of Lebesgue measure $\lvert 4-4R\rvert$. Hence, the case $R=1$ has spectrum of measure zero, and it can be shown to be surely purely singular continuous spectrum. The case $R>1$ is notable for having almost surely pure point spectrum and exhibiting Anderson localization (see Refs. Aubry and André, 1980; Avila and Jitomirskaya, 2006; Avila, 2008).
@@
It will be convenient to set $T_{x}=\mathrm{e}^{\mathrm{i}\nu_{x}}V,T_{y}=\mathrm{e}^{\mathrm{i}\nu_{y}}U$. The group spanned by $T_{x},T_{y}$ is referred to as the group of magnetic translationsZak (1964).
Our goal is now to find an expression for ${\hat{H}}$ in terms of quantum group elements $E,F$, in analogy with the Toda model, in the spirit of Refs. Wiegmann and Zabrodin, 1994; Hatsugai, Kohmoto, and Wu, 1996. Toward this goal, we define a finite-dimensional representation $\varphi$ of $\mathcal{U}_{q}(\mathfrak{sl}_{2})$ in terms of shift and clock operators. The notation suggests that this should be a variant of the representation $\phi$ introduced above in the Toda model setting. Choosing $t=q^{-1}$ in Proposition I.3 and relabeling the Clifford generators as above, we define $\varphi$ as follows:
@@
II Characteristic polynomial and spectral duality
As already mentioned, Hatsuda, Katsura, and TachikawaHatsuda, Katsura, and Tachikawa (2016) found a relationship between the characteristic polynomial of the Hofstadter Hamiltonian $H$, the self-similarity relation defining the fractal properties of the Hofstadter butterfly, and the spectral transformation between the Toda Hamiltonian $H_{\text{Toda}}$ and its modular dual.
Let us denote the polynomial induced by the modular duality between the Toda Hamiltonian $H_{\text{Toda}}$ and its modular dual defined in Ref. Hatsuda, Katsura, and Tachikawa, 2016 by $f=f_{P/Q}(E)$, where $E$ is the energy. We are going to prove that $f$ is the characteristic polynomial $f(E)=\det({\hat{H}}-E)$ of the finite-dimensional Hofstadter Hamiltonian (note that ${\hat{H}}$ depends on $P/Q$, although we suppressed this dependence from the notation).
@@ II Characteristic polynomial and spectral duality
Let us briefly introduce some elements of spectral duality for tridiagonal matrices (see Ref. Molinari, 1997 for more details). The Harper equation can be written in matrix form:
@@
When $\alpha=P/Q$ is a rational number, the energy spectrum of the Hofstadter model splits into $Q$ energy bands. The eigenvalues of ${\hat{H}}$ give us a point in each band, and the rest of the band is obtained by parametrizing over a torus, which coincide with the Brillouin zone, whose coordinates are given by the vector $\vec{\nu}=(\nu_{x},\nu_{y})$. The way $f$ depends on $\vec{\nu}$ is well-understoodChambers (1965); Kohmoto (1989): it is a simple translation as shown by the Chambers relation,
@@
Note that Eq. (17) is in fact a proof of the Chambers relation for the $x$-coordinate. We will see how to obtain the $y$-dependence later after using the representation theory of the quantum group. We will refer to the points $\vec{\nu}=(0,0)$, $\vec{\nu}=(\textstyle\frac{\pi}{2Q},\frac{\pi}{2Q})$, and $\vec{\nu}=(\textstyle\frac{\pi}{Q},\frac{\pi}{Q})$ as the center, mid-band, and corner points, and restrict ourselves to the reduced Brillouin zone $\nu_{x},\nu_{y}\in[0,2\pi/Q]$, except when explicitly noted. This can be done without loss of generality, since the Hamiltonian is periodic in $\nu_{x},\nu_{y}$ with a period $2\pi/Q$ up to unitary transformationsMarra, Citro, and Ortix (2015).
It is well-known that the characteristic polynomial in $f(E,\nu_{x},\nu_{y})$ (of order $Q$) contains only terms with powers $E^{Q-2i}$. This is because odd powers in $E$ necessarily contains terms in $\cos{(2\pi n\alpha-\nu_{y})}$ ), which cancel each other when summed over $n=0,Q-1$ at the mid-band point $\nu_{y}=2\pi/Q$. Consequently, $f(E,\nu_{x},\nu_{y})$ is an even function of $E$ and the spectrum is symmetric under $E\mapsto-E$ if $Q$ is even, while $f(E,\nu_{x},\nu_{y})$ is an odd function and the spectrum at the mid-band point contains $E=0$ if $Q$ is odd, since $f(0,{\pi}/{2Q},{\pi}/{2Q})=0$. It is also well-known that $f(0,{\pi}/{2Q},{\pi}/{2Q})=4(-1)^{Q/2}$ if $Q$ is even, which mandates that $f(0,0,0)=0$ if $Q/2$ is even and $f(0,\pi/Q,\pi/Q)=0$ if $Q/2$ is odd. Consequently, the spectrum contains $E=0$ at the center point if $Q$ is doubly even, at the corner point if $Q$ is singly even, and at the mid-band point if $Q$ is odd, respectively. In particular, the zeros $E=0$ in the spectra when $Q$ is even are doubly degenerate and form Dirac cones with a linear dependence on the momentumWen and Zee (1989).
Let us consider the Hamiltonian $H_{\text{Toda}}$ and its modular dual. Since these commute, the authors in Ref. Hatsuda, Katsura, and Tachikawa, 2016 argue that, by simultaneous diagonalization, we arrive at a pair of difference equations which can be iterated, revealing a tracial relationship of the form
@@
Since $T_{j}$ and $\tilde{T}_{j}$ are defined by shifting along the imaginary axis, at $x=0$ the hyperbolic cosine function becomes a simple cosine function. Then the left-hand side of Eq. (19), evaluated at $x=0$, ought to be the trace of the transfer matrix from Eq. (16) (there is a slight difference due to the convention in Ref. Hatsuda, Katsura, and Tachikawa, 2016 that $[x,p]=2\pi\mathrm{i}\alpha$, but after renormalizing accordingly, the identification with the matrix $T(E)$ in Eq. (16) holds). Then by Theorem II.1 and Eq. (17), the $x$-dependent part in Eq. (19) corresponds to the $\nu_{y}$-term in the Chambers relation.
In particular, since the right-hand side of Eq. (19) is defined via modular duality (swapping $P$ with $Q$), the correction in the definition of $\tilde{T}$ is such that the $x$-dependent terms are always equal for any $P$ and $Q$. Since the polynomial (denoted $P$ in Ref. Hatsuda, Katsura, and Tachikawa, 2016) is defined by effectively setting to zero the $x$-dependent part, we have the identification $P_{\alpha}(E)=(-1)^{Q}f_{P/Q}(E,\pi/2Q,\pi/2Q)$. ∎
Remark II.3.
In Ref. Hatsuda, Katsura, and Tachikawa, 2016 the authors define a polynomial $P$ including the anisotropy parameter $R$, in which case the relevant polynomial is $P_{\alpha}(E)+2R^{Q}$ (note we are adding $2R^{Q}$ rather than $2$).
III Formula for the characteristic polynomial
We now work towards finding a formula for $f(E)=\det({\hat{H}}-E)$. We can exploit the gauge invariance of $H_{\text{Hof}}$ and choose a different form for the vector potential. Following Ref. Wiegmann and Zabrodin, 1994, we can choose $\vec{A}=\alpha/2\cdot(-x-y,\,x+y+1)$. In this gauge, the mid-band point corresponds to $\vec{\nu}=(\pi/2,\pi/2)$. Taking this into account, we can rewrite the representation $\varphi$ from Eq. (14) in terms of magnetic translations:
@@ III Formula for the characteristic polynomial
As a preliminary step, we are interested in the representation $\varphi=\zeta_{0}$ described in Ref. Wiegmann and Zabrodin, 1994, given as
@@
From the representation theory of $\mathcal{U}_{q}(\mathfrak{sl}_{2})$, when $q^{2}$ is a primitive root of unity, we have essentially three families of representations. The first two are low-dimensional, and hence they must be discarded in favor of the last one, which we denote $Z_{a,b}(\lambda)$. This family is topologically parametrized by a $3$-dimensional complex space. However, our system naturally comes with a single parameter $R$, which we have seen in Section I together with the almost Mathieu operator.
@@
By gauge invariance of $H_{\text{Hof}}$, we can compute $\det({\hat{H}}-E)$ when the vector potential is chosen as $\vec{A}=\alpha/2\cdot(-x-y,\,x+y+1)$. In this case, the results in Ref. Wiegmann and Zabrodin, 1994 imply that ${\hat{H}}$ is given as prescribed by Eq. (20) with $\varphi=\zeta_{0}$. According to the classification of the quantum group representations, $\zeta_{0}$ must be equivalent to $Z_{\lambda}$ for some $\lambda$. Indeed, a direct computation (by conjugation with a diagonal matrix) shows that $\lambda=q^{-1}$ (note that, if $\Lambda$ is such diagonal matrix, the system resulting from $\Lambda\zeta_{0}(F)\Lambda^{-1}=\zeta(F)$ is overdetermined). Equivalent representations are related through similar matrices, i.e., which are equivalent up to a similarity (but not necessarily unitary) transformation, and it is well-known that similarity preserves the characteristic polynomial. ∎
Remark III.3.
Note that ${\hat{H}}$ is hermitian, whereas ${\hat{H}}^{\prime}$ is not. However, the matrices ${\hat{H}}$ and ${\hat{H}}^{\prime}$ are equivalent up to a similarity transformation, and therefore isospectral with real spectra: Indeed, ${\hat{H}}^{\prime}$ is pseudo-hermitian, following the definitions introduced by MostafazadehMostafazadeh (2002). Hence, the corresponding Hamiltonian $H^{\prime}$ is an example of a non-Hermitian Hamiltonian with real spectrum which is equivalent up to a similarity transformation to a more conventional hermitian HamiltonianMostafazadeh (2002); Fernández (2015).
Remark III.4.
We see from Definition I.5 that (at least for $\nu_{y}=0$) the diagonal of ${\hat{H}}$ is given by the Chebyshev polynomials of the first kind, in symbols $\theta_{k}=2T_{k-1}(\cos(2\pi\alpha))$. On the other hand, the Hamiltonian ${\hat{H}}^{\prime}$ in the new “quantum group-adjusted” gaugeWiegmann and Zabrodin (1994) shows the Chebyshev polynomials of the second kind on the upper diagonal, that is, ${\hat{H}}^{\prime}_{k,k+1}=U_{k-1}^{2}(\sin(\pi\alpha))$.
Remark III.5.
@@
where in the last equality we reindexed starting from $1$ rather than 00. When $i=0$ we get $x^{N}$ as expected. For the other indices we can sum the two summations and use Lemma III.6 to obtain the result. Analogously, the case where $N$ is even is carried out using the corresponding identities in Lemma III.6. ∎
@@
Combining the previous theorem with Proposition III.2, we get the main result on the polynomial $f$. Recall that $\alpha=P/Q$ is a rational number in reduced form.
Theorem III.9.
@@
Remark III.10.
Let us note that the formula from Theorem III.7 makes at least two aspects clear: if we multiply $y$ by a quantity $x$ and the $b_{i}$’s by the inverse $x^{-1}$, the determinant of $A$ is unaffected. Moreover, when $N$ is odd, the determinant goes to zero if we set $x=0$. We thus recover the well-known result mentioned earlier: the spectrum of the Hofstadter model at the mid-band point always contains 00 when $Q$ is odd.
Remark III.11.
@@
IV Chambers relation and Chebyshev polynomials
@@
Firstly, we know from Section I that these Hamiltonians have the form $(q-q^{-1})\phi(E+F)$ for a suitable representation. Secondly, we obtained an expression of the form ${\hat{H}}^{\prime}=\mathrm{i}(q-q^{-1})\zeta(-E+F)=(q-q^{-1})\zeta(\mathrm{i}^{-1}E+\mathrm{i}F)$. The natural guess away from the mid-band point is to set
@@
We can now use Corollary III.8 to quantify the corner contributions, leading us to the Chambers relation in the coordinates $\kappa_{\pm}$. Let us consider the contribution of one corner, multiplied by $z^{Q}+z^{-Q}$,
@@
We claim that Eq. (25) prescribes the energy dependence on $\vec{\nu}$ (see Theorem IV.2 and its proof for the precise statement).
To prove the claim, inspired by Remark III.4, we could proceed by induction on $Q$ and use the recurrence relations of the Chebyshev polynomials in the proof. Although this seems viable, we will follow a more direct approach.
Lemma IV.1.
@@
can be proved by standard techniques, see for example Ref. Kreft, 1994, Section 4. We will not go into the details here and rather focus on the particular form of the “offset” function $g$.
@@
From Eq. (18) and Proposition III.2 we know we only need to focus on the function $g$. From Eq. (25) and Lemma IV.1, we obtain the corner contribution as follows:
@@
Let us end this section with a small remark about the gauge associated with ${\hat{H}}^{\prime}$ (similar considerations hold for ${\hat{H}}_{c}^{\prime}$). As noted in Ref. Wiegmann and Zabrodin, 1994, the Harper equation at the mid-band point in this setting becomes a difference functional equation:
@@
Recall that the form of the vector potential is chosen to be $\vec{A}=\alpha/2\cdot(-x-y,\,x+y+1)$. Hence a convenient coordinate for the Bloch wave function is $l=n+m$. The Harper equation for $\psi=(\psi_{l})$ is obtained from the equation above by setting $z=q^{l}$ and $\psi_{l}=\Psi(q^{l})$. Notice that $l$ ranges in $1,\dots,2Q$, presenting a “doubling” of the period in comparison with the Harper equation in Section I. Clearly, this is also reflected in the entries of ${\hat{H}}^{\prime}$ where $\pi\alpha$ (rather than $2\pi\alpha$) appears.
@@
More precisely, using Theorems III.7 & III.9, we have that $b_{j}=\sin^{2}(j\pi\alpha)/\sin^{2}(\pi\alpha)$, so that $b_{Q}=0$, and $b_{Q+j}=b_{j}$. Since the variables $b_{j}$ enter the formula only through polynomials which are a modification of the elementary symmetric polynomials, these are computed by expanding the linear factorization of the monic polynomial $\lambda^{2}\prod_{1}^{Q-1}(\lambda-b_{i})^{2}$.